CLASSWORK: A uniform metre rule of mass 90g is pivoted at the 40cm mark. If the metre rule is in equilibrium with an unknown mass, m, placed at the 10cm mark and a 72g mass at the 70cm mark, calculate the value of m.


1
Expert's answer
2020-09-01T10:37:56-0400

Position the ruler so that 0 is on the left, 100 cm is on the right.

Mass of the ruler that hangs on the left and right respectively:


"m_l=m\cdot40/100=0.4m=36\text{ g},\\
m_r=m\cdot(100-40)/100=0.6m_r=54\text{ g}."


Of course, these masses create torque of the ruler, and the lever is a half of the sides around the pivot:


"\tau_l=0.2m_lg=70.56\text{ Nm},\\
\tau_r=0.3m_rg=158.76\text{ Nm}."

Torque created by 72 g mass:


"\tau_1=0.3\cdot72\cdot g=211.68\text{ Nm}."

Torque by the undefined mass:


"\tau=0.3\cdot gm=2.94m."

The ruler is in equilibrium, therefore, torques on the left equal torques on the right:


"70.56+2.94m=158.76+211.68,\\
m=100.6\text{ g}."